At 15 ◦C, the water vapor pressure is 12.79 mmHg. What is the maximum mass of water contained in 2.00·105 L of air at 15 ◦C? The molar mass of water is 18.015 g/mol, 1.00 atm = 760 mmHg, T(K) = T(◦C) + 273.

(A) 49.26 kg (B) 2734 g (C) 2566 g (D) 151.8 g (E) 7.91 g
Answer #1

Given:

P = 12.79 mm Hg

= (12.79/760) atm

= 0.0168 atm

V = 2.00*10^5 L = 200000 L

T = 15.0 oC

= (15.0+273) K

= 288 K

find number of moles using:

P * V = n*R*T

0.0168 atm * 200000 L = n * 0.08206 atm.L/mol.K * 288 K

n = 1.424*10^2 mol

Molar mass of H2O = 18.015 g/mol

use:

mass of H2O,

m = number of mol * molar mass

= 1.424*10^2 mol * 18.015 g/mol

= 2.566*10^3 g

= 2566 g

Answer: C